OpenStudy (anonymous):

The vertex form of the equation of a parabola is y = (x + 3)2 + 53. What is the standard form of the equation? A. y = 6x2 + 9x+ 62 B. y = x2 + 6x + 62 C. y = x2 + 3x + 53 D. y = x2 + 53x + 42

OpenStudy (anonymous):

@caozeyuan

OpenStudy (anonymous):

can u explain pls

OpenStudy (caozeyuan):

notice that the four choices have different linear terms

OpenStudy (anonymous):

what do u mean by linear terms

OpenStudy (caozeyuan):

do you want to do it the proper way or the easy way?

OpenStudy (anonymous):

ik linear is straight but how do u know

OpenStudy (caozeyuan):

y=ax^2=bx+c, bx is the linear term, since it is linear wrt x

OpenStudy (anonymous):

oh ok

OpenStudy (caozeyuan):

sorry, should be y= ax^2+bx+c

OpenStudy (anonymous):

yea, are they always linear?

OpenStudy (caozeyuan):

sometimes this term vanishes, i.e when b=0, for example, y=x^2 is a valid eqn for parabola but it has no linear term

OpenStudy (anonymous):

oh. continue

OpenStudy (caozeyuan):

since we know the four choices have their own linear term, we can compute the linear term only and still have the solution

OpenStudy (anonymous):

C?

OpenStudy (caozeyuan):

if y=(ax+b)^2, the linear term is 2abx

OpenStudy (caozeyuan):

what is the value for a and b in your case?

OpenStudy (anonymous):

im super confused

OpenStudy (caozeyuan):

so expand (x+3)^2, you get x^2+6x+9, agree?

OpenStudy (anonymous):

how did u get this part x^2+6x+9,

OpenStudy (caozeyuan):

(x+3)^2 is (x+3)(x+3)

OpenStudy (caozeyuan):

which is x(x+3)+3(x+3) which is x^2+3x+3x+9, x^2+6x+9